if there are two bodies and we decrease mass of any one body by 9 times. how much should the distance between them be increased to keep force of gravity between them constant.
1
Expert's answer
2020-09-10T14:39:38-0400

According to the Newton's gravity law, the initial force between two masses is the following:


"F_1 = G\dfrac{m_1m_2}{r^2}"

If we increase one of the masses by a factor of 9 ("m_1 \to9m_1" ), then we should increase the distance by a factor of 3 ("r \to 3r") in order to keep the force the same. Let's check:


"F_2 = G\dfrac{(9m_1)m_2}{(3r)^2} = G\dfrac{9m_1m_2}{9r^2} = G\dfrac{m_1m_2}{r^2} = F_1"

Indeen, "F_2 = F_1".

Answer. The distance between masses should be increased by 3 times.